# Self-induced topological edge states in a lattice with onsite nonlinearity

### v1

@@ Abstract
Topological edge states typically arise at the boundaries of topologically nontrivial structures or at interfaces between regions with differing topological invariants. When topological systems are extended into the nonlinear regime, linear topological edge states bifurcate into nonlinear counterparts, and topological gap solitons emerge in the bulk of the structures. Despite extensive studies of these two types of nonlinear states, self-induced topological edge states localized at the physical boundaries of originally nontopological structures remain underexplored. Unlike the previously reported self-induced topological transitions driven by nonlinear couplings, which are conceptually straightforward but less common in realistic interacting systems, here we experimentally realize self-induced topological edge states in a lattice with onsite nonlinearity. Leveraging the strong and tunable nonlinearity of electrical circuits, we systematically investigate the localized states in a nonlinear Su-Schrieffer-Heeger model. Besides revisiting the nonlinear topological edge states and topological gap solitons, we uncover a novel type of self-induced topological edge states which exhibit the hallmark features of linear topological edge states, including sublattice polarization, phase jumps, and decaying tails that approach zero. A distinctive feature of these states is the boundary-induced power threshold for existence. Our results are broadly applicable and can be readily extended to photonic and cold atomic systems, where onsite nonlinearities naturally arise from interparticle interactions. Our work unveils new opportunities for exploring novel correlated topological states of light and matter, and paves the way for the development of robust photonic devices and topological quantum computation.
@@
Topological insulators are physical structures that behave as conventional insulators in the bulk, but conducting on their surfaces due to the existence of topologically protected edge states [1, 2]. The realization of topological insulators has been demonstrated in diverse physical platforms, and the immunity of topological edge states to local deformations and disorders are of great significance for the exciting potential applications [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Extending topological insulators into the nonlinear regime such as by considering the nonlinear response of optical materials under high field intensity [13, 14], the interplay between topology and nonlinearity leads to the formation of the nonlinear topological edge states which bifurcate from the linear edge states and inherit topological protection from the linear counterparts [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In two-dimensional topological systems, due to the modulation instability, the nonlinear edge states develops into the edge solitons where dispersion is balanced with nonlinearity [15, 16, 17, 26, 27, 28, 29, 30]. The bifurcation from the linear topological states to the nonlinear ones is also applicable to the nonlinear higher-order topological systems, where nonlinear corner states are proposed [25, 31].
@@
Figure 1: Schematics of the nonlinear SSH lattices and circuit realization. a, A topological SSH lattice ending with weak bond. Under a weak pump, topological edge states appears at the left edge. b, A structure connected by two SSH lattices with different topological properties. Due to the formation of the domain wall, topological edge state also appears under a weak pump. With the increasing of pump intensity, the topological edge states in both (a) and (b) bifurcate to nonlinear ones. c, For a topological or nontopological SSH lattice, under the external pump, a nonlinearity-induced domain wall appears in the bulk and supports the formation of topological gap solitons. d, A nontopological SSH lattice ending with strong bond. Under the action of onsite nonlinearity, self-induced topological edge states appear at the physical termination. e, Circuit implementation of a unit cell in the lattices shown in (a)-(d). The back-to-back varactor diodes provides the onsite nonlinearity of the SSH model. f, Unit cell of the experimentally realized circuit lattices.
@@
Here, we investigate the topological states in the Su-Schrieffer-Heeger (SSH) model with onsite nonlinearity using electrical circuits [41]. Due to the flexibility of constructing the circuit lattices and the phase-resolved measurement technique, electrical circuits are particularly suited to probe the nonlinear topological physics [42, 43, 44, 45, 46]. As depicted in Fig. 1a, under a weak pump (i.e., in the linear limit), topological edge state appears at the edge of the topological SSH lattice (ending with strong bond). With the increasing of pump intensity, we observe the nonlinear topological edge states bifurcated from the linear counterpart. We experimentally validate the inherited properties of sublattice polarization and phase jump, and find that nonlinearity weakens the localization of the topological edge states. We then implement a structure by connecting two SSH lattices with different topological properties (Fig. 1b). Due to the formation of the domain wall, under a weak pump topological edge state appears as well. We experimentally reveal the similar state properties under increased pump intensity, because the nonlinear topological edge states in Figs. 1a-b have the same physical origin.
@@
Different to the nonlinear topological edge states, topological gap solitons have no direct linear counterparts. For an SSH lattice (Fig. 1c), no matter the lattice is topological or nontopological in the linear limit, under the external pump, a nonlinearity-induced domain wall appears in the bulk and supports the formation of topological gap solitons. We experimentally observe that the left and right tails of the topological gap solitons have opposite sublattice polarizations, and reveal that nonlinearity increases the localization of the topological gap solitons. These features are in contrast to those of nonlinear topological edge states because the physical origins are totally different.
@@
Besides the nonlinear topological edge states and topological gap solitons, we discover both theoretically and experimentally an exotic type of topological states residing at the edge of a semi-infinite lattice ending with strong bond (Fig. 1d). Although the lattice is nontopological in the linear limit, under the action of onsite nonlinearity, it supports the self-induced topological edge states that reside at the physical termination of the structure. The self-induced edge states exhibit the sublattice polarization and phase jump starting from the second site close to the edge. They are the variant of the topological gap solitons under the breaking of the discrete translational symmetry at the structure edge, along with a boundary-induced power threshold for their existence. Specifically, the self-induced topological edge states can be approximately mapped to the linear topological edge state of a semi-infinite SSH lattice (Fig. 1a). Unlike the previously reported self-induced topological transitions driven by nonlinear couplings [42], which are conceptually straightforward but less common in realistic interacting systems, our self-induced topological edge states are realized in a lattice with onsite nonlinearity and feature decaying tails that approach zero, in contrast to the nonlocalized distributions that maintain non-zero plateau levels [42]. Our results are broadly applicable and can be readily extended to photonic and cold atomic systems, where onsite nonlinearities naturally arise from interparticle interactions. Our work unveils new opportunities for exploring novel correlated topological states of light and matter, and paves the way for the development of robust photonic devices and topological quantum computation.
@@ Results
Nonlinear SSH model realized using electrical circuits. All the nonlinear SSH lattices shown in Figs. 1a-d can be realized using electrical circuits. Fig. 1e shows the circuit implementation of a unit cell, where A and B denote the two sublattice sites. For one sublattice, the inductor $L_{\text{g}}$, capacitor $C_{\text{g}}$, and back-to-back varactor diode $C_{\text{v}}$ are wired in parallel. The varactor diode acts as a voltage-dependent variable capacitor with $C_{\text{v}}=C_{\text{L}}+C_{\text{NL}}$, where the linear part $C_{\text{L}}$ is the capacitance at the zero voltage and the nonlinear part can be phenomenologically written as $C_{\text{NL}}=-C_{\text{L}}+\frac{C_{\text{L}}}{\left(1+\left|v/v_{0}\right|\right)^{M}}$ (see Supplementary Note 1). Here, $v_{0}$ and $M$ are constants, and $v$ is the voltage amplitude. The sublattice sites are wired with each other through the intracell coupling capacitor $C_{1}$ and intercell coupling capacitor $C_{2}$.
@@
where $\left[V_{m}^{\text{A}}\left(t\right),V_{m}^{\text{B}}\left(t\right)\right]^{\text{T}}$ are the time-dependent voltages on sites A and B in the $m$th unit cell, $E_{0}={\omega_{0}+}\Delta E$ is equivalent to the constant onsite energy, which includes the resonant frequency of the linear oscillators $\omega_{0}=1/\sqrt{L_{\text{g}}\left(C_{\text{g}}+C_{\text{L}}\right)}$ and the frequency shift induced by the coupling capacitors $\Delta E=-\frac{{C}_{1}+{C}_{2}}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$, $J_{1}=\frac{{C}_{1}}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$ and $J_{2}=\frac{{C}_{2}}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$ are the intracell and intercell coupling coefficients, respectively, and $g\left({V_{m}^{\sigma}}\right)=-\frac{C_{\text{NL}}\left({V_{m}^{\sigma}}\right)}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$ $\left(\sigma=\text{A},\text{B}\right)$ is the voltage-dependent onsite energy. Eq. (12) is valid under $C_{1,2}\ll C_{\text{g}}+C_{\text{L}}$ and $C_{\text{NL}}\ll C_{\text{g}}+C_{\text{L}}$, where the slowly-varying envelope approximation holds (see Supplementary Note 2 for the derivation from the Kirchhoff circcuit equations). Thus, the nonlinear circuit lattice realizes the SSH model with onsite nonlinearity, and the four lattice configurations shown in Figs. 1a-d all can be implemented based on the circuit dimer in Fig. 1e. Fig. 1f shows the unit cell of the experimentally realized circuit lattices.
@@
Nonlinear topological edge states. We first study the nonlinear topological states in the SSH lattices shown in Fig. 1a. The configuration is a typical semi-infinite SSH lattice that ends with a weak bond. In the linear limit, due to the nontrivial topology with the winding number $\mathcal{W}=1$, topological edge state residing at the left edge exists [41]. Under the action of the onsite nonlinearity, by bifurcating from the linear edge state, nonlinear topological edge states with the topological protection inherited from the linear counterpart appear (see Methods for the algorithm and Supplementary Note 3 for the stability analysis). As depicted in Fig. 2a, with the power increasing (defined as the sum of the squares of all the site voltages), the frequency of the nonlinear edge state exhibits a blue shift due to the decreased grounding capacitance, and the state localization becomes weak with a larger participation ratio defined as $(\sum\limits_{m,\sigma}|v_{m}^{\sigma}|^{2})^{2}/\sum\limits_{m,\sigma}|v_{m}^{\sigma}|^{4}$ ($\sigma=\text{A},\text{B}$), particularly when the frequency enters the linear bulk band (denoted by the shaded gray areas). The inset of Fig. 2a shows a typical profile of the nonlinear edge states. Since the nonlinear edge states are bifurcated from the linear counterpart which satisfies $v_{m}^{\text{A}}=\left(-\frac{J_{1}}{J_{2}}\right)^{|m|}v_{0}^{\text{A}}$ and $v_{m}^{\text{B}}=0$, their voltages mainly distribute on sublattice site A with the sublattice pseudospin
@@
We then study the second configuration shown in Fig. 1b. The structure is connected by two lattices with different topological properties defined in the linear limit (winding number $\mathcal{W}=0$ for the left nontopological part and $\mathcal{W}=1$ for the right topological part [47]). Without nonlinearity, a topological edge state is localized at the boundary between the two parts [48, 49]. Similarly, with onsite nonlinearity, nonlinear topological edge states which are bifurcated from the linear edge state appear, as shown in Fig. 2d (see Supplementary Note 5 for the stability analysis). Due to the same physical origin, the nonlinear topological edge states in this lattice also exhibit the sublattice polarization and phase jump, and nonlinearity also weakens the state localization. These properties are confirmed by the experimental measurements presented in in Figs. 2e-f (see Supplementary Note 6 for the experimental implementation, complete voltage spectra, and quantitative analysis). Note that the nonlinear topological edge states are completely different from the nontopological conventional solitons that do not possess these characteristics (see Supplementary Notes 5 for the nontoplogical edge states).
@@
Topological gap solitons. We will now discuss the topological gap solitons in the lattice depicted in Fig. 1b. In contrast to the nonlinear topological edge states, the topological gap solitons lack a direct linear counterpart, although they reside spectrally within the SSH gap. From Fig. 3a, the topological gap solitons bifurcate from the edge of the linear Bloch band and converge to the bulk state in the linear limit (see Supplementary Note 7 for the bifurcation and stability). Under the influence of onsite nonlinearity, a topological interface can be induced at the center unit cell of the lattice [34, 35]. The Dirac mass defined as $m_{\text{Dirac}}=\frac{g\left(V_{m}^{A}\right)-g\left(V_{m}^{B}\right)}{2}$ exhibits an inversion ($m_{\text{Dirac}}>0$ for the left part and $m_{\text{Dirac}}<0$ for the right part), and the topological gap solitons emerge as the Jackiw-Rebbi-type Dirac boundary modes (see Supplementary Note 7 for the physical interpretation). The nonlinearity-induced interface can also be interpreted as an impurity potential that splits the the original lattice into two topological regions, and the topological gap soliton represents a combination of the two topological edge states [20].
@@
Self-induced topological edge states. The nonlinear topological states we currently implement rely on a topological edge, whether it is the physical termination of a lattice or a topological interface induced by nonlinearity. Hereafter, we turn to a semi-infinite lattice ending with a strong bond (Fig. 1c), and show the emergence of an exotic type of nonlinear topological states that reside at the edge of a nonlinear nontopological lattice.
@@ Methods
Sample fabrication and measurement. The circuit parameters in the four SSH circuits are $C_{\text{g}}=4.7~{}\text{nF}$, $L_{\text{g}}=15~{}\mu\text{H}$, $C_{\text{L}}=73.48~{}\text{pF}$, $v_{0}=2.1935$, and $M=0.4548$. For the lattice shown in Fig. 1a, the intracell and intercell coupling capacitors are $C_{1}=180~{}\text{pF}$ and $C_{2}=560~{}\text{pF}$, respectively. For the lattice shown in Fig. 1b, the topological part has $C_{1}=180~{}\text{pF}$ and $C_{2}=560~{}\text{pF}$, while the nontopological part has $C_{1}=560~{}\text{pF}$ and $C_{2}=180~{}\text{pF}$. An additional capacitance $C_{2}-C_{1}$ is added to the grounding capacitance of the interface circuit node. For the lattices shown in Fig. 1c-d, the intracell and intercell coupling capacitors are $C_{1}=560~{}\text{pF}$ and $C_{2}=180~{}\text{pF}$, respectively.

### v2

@@ Abstract
Topological edge states typically arise at the boundaries of topologically nontrivial structures or at interfaces between regions with differing topological invariants. When topological systems are extended into the nonlinear regime, linear topological edge states bifurcate into nonlinear counterparts, and topological gap solitons emerge in the bulk of the structures. Extensive studies of these two types of nonlinear states have been carried out, and self-induced topological edge states localized at the physical boundaries of originally nontopological structures have also been discussed. Unlike the previously reported self-induced topological transitions driven by nonlinear couplings, which are conceptually straightforward but less common in realistic interacting systems, here we experimentally realize self-induced topological edge states in a lattice with onsite nonlinearity. Leveraging the strong and tunable nonlinearity of electrical circuits, we systematically investigate the localized states in a nonlinear Su-Schrieffer-Heeger model. Besides revisiting the nonlinear topological edge states and topological gap solitons, we uncover a novel type of self-induced topological edge states which exhibit the hallmark features of linear topological edge states, including sublattice polarization, phase jumps, and decaying tails that approach zero. A distinctive feature of these states is the boundary-induced power threshold for existence. Our results are broadly applicable and can be readily extended to photonic and cold atomic systems, where onsite nonlinearities naturally arise from interparticle interactions. Our work unveils new opportunities for exploring novel correlated topological states of light and matter, and paves the way for the development of robust photonic devices and topological quantum computation.
@@
Topological insulators are physical structures that behave as conventional insulators in the bulk, but conducting on their surfaces due to the existence of topologically protected edge states [1, 2]. The realization of topological insulators has been demonstrated in diverse physical platforms, and the immunity of topological edge states to local deformations and disorders are of great significance for the exciting potential applications [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Extending topological insulators into the nonlinear regime such as by considering the nonlinear response of optical materials under high field intensity [13, 14], the interplay between topology and nonlinearity leads to the formation of the nonlinear topological edge states which bifurcate from the linear edge states and inherit topological protection from the linear counterparts [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In two-dimensional topological systems, due to the modulation instability, the nonlinear edge states develops into the edge solitons where dispersion is balanced with nonlinearity [15, 16, 17, 28, 29, 30, 31, 32]. The bifurcation from the linear topological states to the nonlinear ones is also applicable to the nonlinear higher-order topological systems, where nonlinear corner states are proposed [25, 33, 34].
@@
Figure 1: Schematics of the nonlinear SSH circuit. a, Circuit implementation of the SSH lattices. The back-to-back varactor diodes provides the onsite nonlinearity of the SSH model. b, Unit cell of the experimentally realized circuit lattices.
@@
Following the implementation in an optical platform [42], here, we investigate the topological states in the Su-Schrieffer-Heeger (SSH) model with onsite nonlinearity using electrical circuits [45]. Due to the flexibility of constructing the circuit lattices and the phase-resolved measurement technique, electrical circuits are particularly suited to probe the nonlinear topological physics [46, 47, 48, 49, 50]. Under a weak pump (i.e., in the linear limit), topological edge state appears at the edge of the topological SSH lattice (ending with strong bond). With the increasing of pump intensity, we observe the nonlinear topological edge states bifurcated from the linear counterpart. We experimentally validate the inherited properties of sublattice polarization and phase jump, and find that nonlinearity weakens the localization of the topological edge states. We then implement a structure by connecting two SSH lattices with different topological properties. Due to the formation of the domain wall, under a weak pump topological edge state appears as well. We experimentally reveal the similar state properties under increased pump intensity, because the nonlinear topological edge states have the same physical origin.
@@
Different to the nonlinear topological edge states, topological gap solitons have no direct linear counterparts. For an SSH lattice, no matter the lattice is topological or nontopological in the linear limit, under the external pump, a nonlinearity-induced domain wall appears in the bulk and supports the formation of topological gap solitons. We experimentally observe that the left and right tails of the topological gap solitons have opposite sublattice polarizations, and reveal that nonlinearity increases the localization of the topological gap solitons. These features are in contrast to those of nonlinear topological edge states because the physical origins are totally different.
@@
Besides the nonlinear topological edge states and topological gap solitons, we discover both theoretically and experimentally an exotic type of topological states residing at the edge of a semi-infinite lattice ending with strong bond. Although the lattice is nontopological in the linear limit, under the action of onsite nonlinearity, it supports the self-induced topological edge states that reside at the physical termination of the structure [51]. The self-induced edge states exhibit the sublattice polarization and phase jump starting from the second site close to the edge. They are the variant of the topological gap solitons under the breaking of the discrete translational symmetry at the structure edge, along with a boundary-induced power threshold for their existence. Specifically, the self-induced topological edge states can be approximately mapped to the linear topological edge state of a semi-infinite SSH lattice. Unlike the previously reported self-induced topological transitions driven by nonlinear couplings [46], which are conceptually straightforward but less common in realistic interacting systems, our self-induced topological edge states are realized in a lattice with onsite nonlinearity and feature decaying tails that approach zero, in contrast to the nonlocalized distributions that maintain non-zero plateau levels [46]. Our results are broadly applicable and can be readily extended to photonic and cold atomic systems, where onsite nonlinearities naturally arise from interparticle interactions. Our work unveils new opportunities for exploring novel correlated topological states of light and matter, and paves the way for the development of robust photonic devices and topological quantum computation.
@@ Results
Nonlinear SSH model realized using electrical circuits. All the nonlinear SSH lattices can be realized using electrical circuits. Fig. 1a shows the circuit implementation of a unit cell, where A and B denote the two sublattice sites. For one sublattice, the inductor $L_{\text{g}}$, capacitor $C_{\text{g}}$, and back-to-back varactor diode $C_{\text{v}}$ are wired in parallel. The varactor diode acts as a voltage-dependent variable capacitor with $C_{\text{v}}=C_{\text{L}}+C_{\text{NL}}$, where the linear part $C_{\text{L}}$ is the capacitance at the zero voltage and the nonlinear part can be phenomenologically written as $C_{\text{NL}}=-C_{\text{L}}+\frac{C_{\text{L}}}{\left(1+\left|v/v_{0}\right|\right)^{M}}$ (see Supplementary Note 1). Here, $v_{0}$ and $M$ are constants, and $v$ is the voltage amplitude. The sublattice sites are wired with each other through the intracell coupling capacitor $C_{1}$ and intercell coupling capacitor $C_{2}$.
@@
where $\left[V_{m}^{\text{A}}\left(t\right),V_{m}^{\text{B}}\left(t\right)\right]^{\text{T}}$ are the time-dependent voltages on sites A and B in the $m$th unit cell, $E_{0}={\omega_{0}+}\Delta E$ is equivalent to the constant onsite energy, which includes the resonant frequency of the linear oscillators $\omega_{0}=1/\sqrt{L_{\text{g}}\left(C_{\text{g}}+C_{\text{L}}\right)}$ and the frequency shift induced by the coupling capacitors $\Delta E=-\frac{{C}_{1}+{C}_{2}}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$, $J_{1}=\frac{{C}_{1}}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$ and $J_{2}=\frac{{C}_{2}}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$ are the intracell and intercell coupling coefficients, respectively, and $g\left({V_{m}^{\sigma}}\right)=-\frac{C_{\text{NL}}\left({V_{m}^{\sigma}}\right)}{2\left(C_{\text{g}}+C_{\text{L}}\right)}{\omega_{0}}$ $\left(\sigma=\text{A},\text{B}\right)$ is the voltage-dependent onsite energy. Eq. (12) is valid under $C_{1,2}\ll C_{\text{g}}+C_{\text{L}}$ and $C_{\text{NL}}\ll C_{\text{g}}+C_{\text{L}}$, where the slowly-varying envelope approximation holds (see Supplementary Note 2 for the derivation from the Kirchhoff circcuit equations). Thus, the nonlinear circuit lattice realizes the SSH model with onsite nonlinearity, and the four lattice configurations all can be implemented based on the circuit dimer in Fig. 1a. Fig. 1b shows the unit cell of the experimentally realized circuit lattices.
@@
Nonlinear topological edge states. We first study the nonlinear topological states in the first SSH lattice. The configuration is a typical semi-infinite SSH lattice that ends with a weak bond. In the linear limit, due to the nontrivial topology with the winding number $\mathcal{W}=1$, topological edge state residing at the left edge exists [45]. Under the action of the onsite nonlinearity, by bifurcating from the linear edge state, nonlinear topological edge states with the topological protection inherited from the linear counterpart appear (see Methods for the algorithm and Supplementary Note 3 for the stability analysis). As depicted in Fig. 2a, with the power increasing (defined as the sum of the squares of all the site voltages), the frequency of the nonlinear edge state exhibits a blue shift due to the decreased grounding capacitance, and the state localization becomes weak with a larger participation ratio defined as $(\sum\limits_{m,\sigma}|v_{m}^{\sigma}|^{2})^{2}/\sum\limits_{m,\sigma}|v_{m}^{\sigma}|^{4}$ ($\sigma=\text{A},\text{B}$), particularly when the frequency enters the linear bulk band (denoted by the shaded gray areas). The inset of Fig. 2a shows a typical profile of the nonlinear edge states. Since the nonlinear edge states are bifurcated from the linear counterpart which satisfies $v_{m}^{\text{A}}=\left(-\frac{J_{1}}{J_{2}}\right)^{|m|}v_{0}^{\text{A}}$ and $v_{m}^{\text{B}}=0$, their voltages mainly distribute on sublattice site A with the sublattice pseudospin
@@
We then study the second configuration. The structure is connected by two lattices with different topological properties defined in the linear limit (winding number $\mathcal{W}=0$ for the left nontopological part and $\mathcal{W}=1$ for the right topological part [52]). Without nonlinearity, a topological edge state is localized at the boundary between the two parts [53, 54]. Similarly, with onsite nonlinearity, nonlinear topological edge states which are bifurcated from the linear edge state appear, as shown in Fig. 2d (see Supplementary Note 5 for the stability analysis). Due to the same physical origin, the nonlinear topological edge states in this lattice also exhibit the sublattice polarization and phase jump, and nonlinearity also weakens the state localization. These properties are confirmed by the experimental measurements presented in in Figs. 2e-f (see Supplementary Note 6 for the experimental implementation, complete voltage spectra, and quantitative analysis). Note that the nonlinear topological edge states are completely different from the nontopological conventional solitons that do not possess these characteristics (see Supplementary Notes 5 for the nontoplogical edge states).
@@
Topological gap solitons. We will now discuss the topological gap solitons in the third kind of the lattice. In contrast to the nonlinear topological edge states, the topological gap solitons lack a direct linear counterpart, although they reside spectrally within the SSH gap. From Fig. 3a, the topological gap solitons bifurcate from the edge of the linear Bloch band and converge to the bulk state in the linear limit (see Supplementary Note 7 for the bifurcation and stability). Under the influence of onsite nonlinearity, a topological interface can be induced at the center unit cell of the lattice [37, 38]. The Dirac mass defined as $m_{\text{Dirac}}=\frac{g\left(V_{m}^{A}\right)-g\left(V_{m}^{B}\right)}{2}$ exhibits an inversion ($m_{\text{Dirac}}>0$ for the left part and $m_{\text{Dirac}}<0$ for the right part), and the topological gap solitons emerge as the Jackiw-Rebbi-type Dirac boundary modes (see Supplementary Note 7 for the physical interpretation). The nonlinearity-induced interface can also be interpreted as an impurity potential that splits the the original lattice into two topological regions, and the topological gap soliton represents a combination of the two topological edge states [20].
@@
Self-induced topological edge states. The nonlinear topological states we currently implement rely on a topological edge, whether it is the physical termination of a lattice or a topological interface induced by nonlinearity. Hereafter, we turn to a semi-infinite lattice ending with a strong bond, and show the emergence of an exotic type of nonlinear topological states that reside at the edge of a nonlinear nontopological lattice.
@@ Methods
Sample fabrication and measurement. The circuit parameters in the four SSH circuits are $C_{\text{g}}=4.7~{}\text{nF}$, $L_{\text{g}}=15~{}\mu\text{H}$, $C_{\text{L}}=73.48~{}\text{pF}$, $v_{0}=2.1935$, and $M=0.4548$. For the first lattice, the intracell and intercell coupling capacitors are $C_{1}=180~{}\text{pF}$ and $C_{2}=560~{}\text{pF}$, respectively. For the second lattice, the topological part has $C_{1}=180~{}\text{pF}$ and $C_{2}=560~{}\text{pF}$, while the nontopological part has $C_{1}=560~{}\text{pF}$ and $C_{2}=180~{}\text{pF}$. An additional capacitance $C_{2}-C_{1}$ is added to the grounding capacitance of the interface circuit node. For the third and fourth lattices, the intracell and intercell coupling capacitors are $C_{1}=560~{}\text{pF}$ and $C_{2}=180~{}\text{pF}$, respectively.